# Modeling Metabolic State Transitions in Obesity Using a Time-Varying Lambda–Omega Framework

### 06 Mar 2026 v1

@@ Abstract
Obesity does not emerge abruptly; rather, it develops gradually over extended periods. The gradual progression often prevents early recognition of physiological changes until excess adiposity is established. A common belief is that weight reduction can be achieved simply by ”eating less and moving more”. Although reductions in caloric intake and increases in physical activity are fundamental principles of weight management, this perspective oversimplifies a complex and adaptive biological system. Metabolic rate, hormonal regulation, behavioral factors, and compensatory physiological responses all influence the body’s resistance to changes in weight. During weight loss, reduced metabolic rate and increased efficiency make maintaining a caloric deficit increasingly difficult. Conversely, during periods of overfeeding, resting metabolic rate, the thermic effect of food, and non-exercise activity thermogenesis increase with rising body weight, partially offsetting the caloric surplus and slowing weight gain. However, these compensatory responses are asymmetrical, with stronger and more persistent adaptations to underfeeding than to overfeeding. This asymmetry helps explain why weight gain often occurs gradually and why sustained weight loss is biologically challenging. In this work, we employ a lambda-omega model from dynamical systems theory to describe metabolic regulation in response to lifestyle perturbations. We introduce time-varying parameters that allow the regulatory coefficients to evolve gradually under sustained environmental and physiological stressors. By allowing lambda(t) and omega(t) to vary over time, the model captures progressive shifts in the metabolic set-point and deformation of the underlying dynamical landscape. This framework enables exploration of transitions between metabolic states and long-term adaptations that shape trajectories of weight gain and loss.
@@ 1 Introduction
It was subsequently adopted in the mathematical biology literature as a canonical model for limit-cycle oscillations and wave propagation (e.g., Murray Murray (2002)). Further theoretical developments, such as those by Sherratt Sherratt (1994), explored the formation of wave trains and spatial patterns arising from $\lambda$$\omega$ kinetics, while more recent work (e.g., Flach Flach (2007)) has continued to use this formulation as a baseline model for nonlinear oscillatory phenomena.

### 17 Mar 2026 v2

@@ Abstract
Obesity develops gradually over extended periods rather than emerging abruptly. Because this progression is slow and often subtle, physiological changes may remain unrecognized until excess adiposity is already established. Weight reduction is commonly framed as a simple consequence of “eating less and moving more”; however, this view overlooks the complex and adaptive nature of human metabolic regulation. During weight loss, resting metabolic rate often decreases and metabolic efficiency increases, making it progressively more difficult to maintain a caloric deficit. In contrast, during periods of overfeeding, increases in resting metabolic rate, thermogenesis, and spontaneous physical activity partially counteract excess energy intake and slow weight gain. Importantly, these compensatory responses are asymmetric, with stronger and more persistent adaptations occurring during underfeeding than during overfeeding. This asymmetry contributes to the gradual development of obesity and the biological difficulty of sustaining long-term weight loss. Here we introduce a dynamical-systems framework based on a lambda–omega model to describe metabolic regulation under lifestyle perturbations. By allowing the regulatory parameters to vary over time, the model captures progressive shifts in metabolic set-points and enables exploration of transitions between metabolic states underlying long-term weight gain and loss trajectories.
@@ 1 Introduction
It was subsequently adopted in the mathematical biology literature as a canonical model for limit-cycle oscillations and wave propagation (e.g., Murray Murray (2002), Lederman Lederman et al. (2022)). Further theoretical developments, such as those by Sherratt Sherratt (1994), explored the formation of wave trains and spatial patterns arising from $\lambda$$\omega$ kinetics, while more recent work (e.g., Flach Flach (2007)) has continued to use this formulation as a baseline model for nonlinear oscillatory phenomena.